# thanks: Corresponding author

### v2

@@ V Conclusion
In this work, we introduced the concept of mutually equi-biased bases (MEBs), a generalization of the mutually unbiased bases (MUBs). For such bases, we relaxed the condition of mutual maximum unpredictability that holds for MUBs; instead, we associated a probability distribution $q=(q_{0},\ldots,q_{d-1})$ with each of the MEBs. This distribution captures the predictability of the outcomes of a measurement in one basis when the system is prepared in a state from the other basis. However, the information extracted from the measurement is unbiased, meaning that it cannot reveal which state from which basis the system is prepared in, and which basis is used for the measurement.
We derived a complete set of $d+1$ MEBs for $d=2$ and $d=3$, and found that although for $d=2$ a complete set of three MEBs exist for all values of the index $\mu=\sum_{k=0}^{d-1}q_{k}^{2}$, for $d\geq 3$ this is not the case. More precisely, for $d=3$, the constraints imposed by the mutual equi-biasedness limits the range of the index $\mu$ to the feasible region $1/3\leq\mu\leq 1/2$ within which a set of four MEBs exists. Furthermore, in addition to this restriction on $\mu$, for a given $\mu$ in the feasible region the components of the probability distribution $q=(q_{0},q_{1},q_{2})$ are also limited to some specific discrete points; only one probability distribution qq—up to permutation of its components—can generate a complete set of four MEBs.
@@
We examined our entanglement witness for $d\times d$ isotropic states, and, in the case $d=3$, we evaluated it using a set of $L$ MEBs, and presented the results through illustrative plots for $L=2,3,4$. We showed that in the absence of a complete set of MUBs, having a set of MEBs with at least one additional basis can already enhance the construction of finer entanglement witnesses.
These results provide a deeper understanding of the structure and properties of MEBs, offering new tools for entanglement detection and quantum measurement theory. The challenging dimension $d=6$ is currently under investigation to determine whether there exist values of the index $\mu$ for which one can construct a set of at least four MEBs—exceeding a set of three MUBs currently known for this dimension. We hope that this work will serve as a stepping stone for further research, ultimately advancing our understanding of MUBs. In particular, for dimensions where no more than three MUBs are known, even a set of four MEBs with $\mu>1/d$ could shed light on the challenging open problem of constructing complete set of MUBs.
@@ acknowledgment
This work is based upon research funded by Iran National Science Foundation (INSF) under project No. 4021765.
Appendix A Derivation of the eigenvalues of $G$
@@
where ${\mathbb{I}}_{d-1}$ is the unit matrix of dimension (d1)(d-1), and the $(d-1)\times(d-1)$ matrix $N_{d-1}$, defined by Eq. (52), forms the blocks of the Gram matrix $G$. Note here that, to maintain clarity, we explicitly denote the size of each submatrix with a subscript. This simple block form of $G$ allows us to provide the following direct product representation
@@
where $\lambda_{i}=|N_{i,d-i}|$, and $\lfloor\frac{d-1}{2}\rfloor$ denotes the integral part of $\frac{d-1}{2}$, i.e., $\lfloor\frac{d-1}{2}\rfloor=\frac{d-1}{2}$ if $d$ is odd, and $\lfloor\frac{d-1}{2}\rfloor=\frac{d}{2}-1$ if $d$ is even. Note that all d1d-1 eigenvalues of $N_{d-1}$ occur in positive-negative pairs, except for the last one, which appears alone. Obviously, this lone eigenvalue, i.e., $\lambda_{\frac{d}{2}}$, exists only when $d$ is even.
@@
Eigenvalues of $G$ when $d$ is odd.—In this case, for $i\in\{1,\ldots,(d-1)/2\}$, the nondegenerate eigenvalues are
@@
Eigenvalues of $G$ when $d$ is even.—In this case, the eigenvalues are the same as above, here for $i\in\{1,\ldots,d/2-1\}$, along with the two new eigenvalues $1+(L-1)|N_{d/2,d/2}|$ and $1-|N_{d/2,d/2}|$, which have multiplicities $1$ and L1L-1, respectively.

### v3

@@ V Conclusion
In this work we introduced the concept of mutually equibiased bases, a generalization of the mutually unbiased bases. For such bases, we relaxed the condition of mutual maximum unpredictability that holds for MUBs; instead, we associated a probability distribution $q=(q_{0},\ldots,q_{d-1})$ with each of the MEBs. This distribution captures the predictability of the outcomes of a measurement in one basis when the system is prepared in a state from the other basis. However, the information extracted from the measurement is unbiased, meaning that it cannot reveal which state from which basis the system is prepared in and which basis is used for the measurement.
We derived a set of $d+1$ MEBs for $d=2$ and $d=3$ and found that although for $d=2$ a complete set of three MEBs exist for all values of the index $\mu=\sum_{k=0}^{d-1}q_{k}^{2}$, for $d\geq 3$ this is not the case. More precisely, for $d=3$, the constraints imposed by the mutual equibiasedness limits the range of the index $\mu$ to the feasible region $1/3\leq\mu\leq 1/2$ within which a set of four MEBs exists. Furthermore, in addition to this restriction on $\mu$, for a given $\mu$ in the feasible region the components of the probability distribution $q=(q_{0},q_{1},q_{2})$ are also limited to some specific discrete points; only one probability distribution $q$, up to permutation of its components, can generate a set of four MEBs.
@@
We examined our entanglement witness for $d\times d$ isotropic states and, in the case $d=3$, we evaluated it using a set of $L$ MEBs; we presented the results through illustrative plots for $L=2,3,4$. We showed that in the absence of a complete set of MUBs, having a set of MEBs with at least one additional basis can enhance the construction of finer entanglement witnesses.
These results provide a deeper understanding of the structure and properties of MEBs, offering new tools for entanglement detection and quantum measurement theory. The challenging dimension $d=6$ is currently under investigation to determine whether there exist values of the index $\mu$ for which one can construct a set of at least four MEBs, exceeding a set of three MUBs currently known for this dimension. We hope that this work will serve as a stepping stone for further research, ultimately advancing our understanding of MUBs. In particular, for dimensions where no more than three MUBs are known, even a set of four MEBs with $\mu>1/d$ could shed light on the challenging open problem of constructing complete set of MUBs.
@@ acknowledgment
This work was based upon research funded by Iran National Science Foundation under Project No. 4021765.
Appendix A Proof of Proposition 2
Proof.
Having the first unitary matrix $\mathcal{U}$, we have to find two other unitary matrices $\mathcal{V}$ and $\mathcal{W}$ with the same moduli of matrix elements as $\mathcal{U}$. The second unitary matrix $\mathcal{V}$ can be constructed as $\mathcal{V}=\mathcal{D}_{\alpha}\mathcal{U}$ where $\mathcal{D}_{\alpha}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\alpha_{1}},{\mathrm{e}}^{i\alpha_{2}}\}$. Regarding the probability $q=(q_{0},q_{1},q_{2})$, although both $\mathcal{U}$ and $\mathcal{V}$ generate bases that are equibiased with respect to the standard basis, they are not necessarily equibiased with respect to each other. For this to hold, $\mathcal{U}$ and $\mathcal{V}$ must be mutually equibiased, i.e., the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$ must generate the same probability matrix $\mathcal{Q}$, up to a permutation of its entries. This places some constraints on phases $\alpha_{1}$ and $\alpha_{2}$.
Here $\sigma\in S_{3}$ denotes a permutation of three elements $\{0,1,2\}$, where $S_{3}$ is the symmetric group of degree $3$. For clarity, the six operations of $S_{3}$ can be written as the identity $\mathrm{id}$, the transpositions $(01)$, $(02)$, $(12)$, and the forward and backward cycles $(012)$ and $(210)$. After some calculations, we get
which holds for $\mu\in[1/3,5/9]$.
Equation (70) imposes a constraint on the allowed values of the phases $\alpha_{1}$ and $\alpha_{2}$ under which two unitary matrices $\mathcal{U}$ and $\mathcal{V}$ are mutually equibiased. A contour plot of this equation in $\alpha_{1}\alpha_{2}$-plane for different values of $\mu$ is illustrated in Fig. 4. The contours that are bounded within the hexagon correspond to $\mu\in[1/3,1/2]$ and those outside the hexagon, i.e., within the two outer triangles, correspond to $\mu\in(1/2,5/9]$.
Figure 4: Contour plot of the left-hand side of Eq. (70) in $\alpha_{1}\alpha_{2}$ plane for $\alpha_{1},\alpha_{2}\in(-\pi,+\pi)$. (a) The inner boundary (green ellipse) corresponds to $\mu=1/3$, (b) the outer boundary (red hexagon) corresponds to $\mu=1/2$, and (c) two little pink contours are plotted for $\mu\approx 5/9$. The exact value for $\mu=5/9$ correspond to two points at $(2\pi/3,-2\pi/3)$ and $(-2\pi/3,2\pi/3)$. The $\alpha_{1}\alpha_{2}$ plane is divided by dashed grid lines to make it easier to identify the quadrant of the angles $\alpha_{1}$ and $\alpha_{2}$.
Clearly, the constraint imposed by Eq. (70) is invariant under the symmetric group $S_{3}$; it depends only on $\mu$ but not on where the probability $q=(q_{0},q_{1},q_{2})$ is located in the feasible region. However, a look at the left-hand sides of Eqs. (67)-(69) reveals that this set is invariant under a subgroup of $S_{3}$, namely, the cyclic group $C_{3}=\{\mathrm{id},(012),(210)\}$. The set, however, will possess the full symmetry under $S_{3}$ if and only if two out of the three terms $\cos{\alpha_{1}}$, $\cos{\alpha_{2}}$, and $\cos{\alpha_{12}}$ become equal. This results in three possible cases as
  1. (a) :
    $\alpha_{1}=0,\quad\quad\alpha_{2}=\theta^{\mu}$,
  2. (b) :
    $\alpha_{1}=\theta^{\mu},\quad\;\alpha_{2}=0$,
  3. (c) :
    $\alpha_{1}=\theta^{\mu},\quad\;\alpha_{2}=\theta^{\mu}$,
where, from Eq. (70), $\theta^{\mu}$ is obtained as Eq. (28). The index condition $\mu\leq 1/2$ is obtained from $|\cos{\theta^{\mu}}|\leq 1$. Using this in Eqs. (67)-(69), we get for case (a) the set of equations
Depending on the choice of the permutation $\sigma\in S_{3}$, the set of equations determines the probability distribution $q=(q_{0},q_{1},q_{2})$. More precisely, the choice $\sigma=\mathrm{id}$ fixes $q_{1}=q^{\mu}_{0}$, so that $q=(q^{\mu}_{\pm},q^{\mu}_{0},q^{\mu}_{\mp})$, where $q^{\mu}_{0}=(1-\mu)/2$ and $q^{\mu}_{\pm}$ is given in Eq. (30). Similarly, the cycle $\sigma=(012)$ fixes $q_{0}=q^{\mu}_{0}$, giving $q=(q^{\mu}_{0},q^{\mu}_{\pm},q^{\mu}_{\mp})$, while the cycle $\sigma=(210)$ yields $q=(q^{\mu}_{\pm},q^{\mu}_{\mp},q^{\mu}_{0})$. In the case of the transpositions $(01)$, $(02)$, and $(12)$, the set possesses solution only for $\mu=1/3$, corresponding to the uniform distribution $q=(1/3,1/3,1/3)$.
In a similar manner, for cases (b) and (c), the set of equations (67)-(69) reduce to
and
respectively. In each case, the corresponding set of equations gives, up to a permutation, the same set of solutions (29) and (30) for the probabilities $(q_{0},q_{1},q_{2})$. Irrespective of which case (a), (b), or (c) is considered, only three out of six possible permutations $\sigma$, namely, those corresponding to the cyclic group $C_{3}=\{\mathrm{id},(012),(210)\}$, are contributed in the solutions. Table 1 presents all possible solutions of Eqs. (67)-(69), corresponding to the three cases (a), (b), and (c) and three permutations $\mathrm{id}$, $(012)$, and $(210)$.
Table 1: Possible solutions of Eqs. (67)-(69).
$\sigma=\mathrm{id}$ $\sigma=(012)$ $\sigma=(210)$
To complete the characterization of the set of MEBs, we have to construct the third unitary matrix $\mathcal{W}$ as $\mathcal{W}=\mathcal{D}_{\beta}\mathcal{U}$ where $\mathcal{D}_{\beta}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\beta_{1}},{\mathrm{e}}^{i\beta_{2}}\}$. Similar to the arguments given above, we require that the unitaries $\mathcal{W}\mathcal{U}$ and $\mathcal{W}\mathcal{V}$ generate the unistochastic matrix $\mathcal{Q}$, up to a permutation of its rows and columns. The first requirement is achieved if we select the pair $(\beta_{1},\beta_{2})$ from case (a), (b), or (c), mentioned above for $(\alpha_{1},\alpha_{2})$. To ensure that $\mathcal{W}\mathcal{V}$ also generates the unistochastic matrix $\mathcal{Q}$, we have to choose the pair $(\beta_{1},\beta_{2})$ distinct from the pair $(\alpha_{1},\alpha_{2})$. ∎
As we have seen, the mutual equibiasedness imposes some constraints on the index $\mu=q_{0}^{2}+q_{1}^{2}+q_{2}^{2}$ of the distribution $q=(q_{0},q_{1},q_{2})$, meaning that, not any value of $\mu$ is allowed to generate a bistochastic matrix from the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$. The following remark clarifies how the feasible region is derived.
Remark 7.
The feasible region within the probability triangle is determined by applying a sequence of constraints arising from physical considerations. These conditions successively restrict the parameter space, resulting in a subset that satisfies all necessary properties. Figure 5 outlines how the feasible region is constructed in this case, with each panel depicting a specific constraint imposed on the probability triangle, as detailed below.
Figure 5(a) shows a subset of the probability triangle that satisfies the chain-link conditions (23). Corresponding to each probability distribution $q=(q_{0},q_{1},q_{2})$ within the subset, the associated unitary matrix $\mathcal{U}$ generates the corresponding unistochastic matrix $\mathcal{Q}$.
For $\mu\in[1/3,1)$, the right-hand side of Eq. (70) can take any nonpositive value $(-\infty,0]$, but the left-hand side is restricted to the interval [3/2,3][-3/2,3]. This poses restrictions on the index $\mu$ in the sense that only distributions with index $\mu\in[1/3,5/9]$ can generate unistochastic matrices from $\mathcal{V}^{\dagger}\mathcal{U}$. The resulting feasible region is depicted in Fig. 5(b).
Equations (67)-(69) pose some chain-link conditions on the four segments $q_{0}$, $q_{1}$, $q_{2}$, and $\sqrt{q_{i}}$ as well, i.e., the nontrivial conditions $q_{i}\leq q_{j}+q_{k}+\sqrt{q_{j}}$ hold, for $(i,j,k)$ taken from $(0,1,2)$. After imposing these constraints, the resulting feasible domain is reduced to the region shown in Fig. 5(c).
Finally, Eq. (28), together with the condition $|\cos{\theta^{\mu}}|\leq 1$, leads to the constraint $\mu\leq 1/2$. Figure 5(d) illustrates the feasible region corresponding to $\mu\in[1/3,1/2]$, consistent with all previously imposed constraints.
Figure 5: Geometric representation of the probability distribution $q=(q_{0},q_{1},q_{2})$ and the construction of the feasible region. (a) Equilateral triangle associated with the probability distribution $q=(q_{0},q_{1},q_{2})$, with its unistochastic subset. All plots are in the $\delta_{2}\delta_{1}$ plane. Each panel corresponds to a constraint applied in succession. See Remark 7 for details.
Appendix B Derivation of the eigenvalues of $G$
@@
where ${\mathbb{I}}_{d-1}$ is the unit matrix of dimension (d1)(d-1) and the $(d-1)\times(d-1)$ matrix $N_{d-1}$, defined by Eq. (44), forms the blocks of the Gram matrix $G$. Note here that, to maintain clarity, we explicitly denote the size of each submatrix with a subscript. This simple block form of $G$ allows us to provide the direct product representation
@@
where $\lambda_{i}=|N_{i,d-i}|$ and $\lfloor\frac{d-1}{2}\rfloor$ denotes the integral part of $\frac{d-1}{2}$, i.e., $\lfloor\frac{d-1}{2}\rfloor=\frac{d-1}{2}$ if $d$ is odd, and $\lfloor\frac{d-1}{2}\rfloor=\frac{d}{2}-1$ if $d$ is even. Note that all d1d-1 eigenvalues of $N_{d-1}$ occur in positive-negative pairs, except for the last one, which appears alone. Obviously, this lone eigenvalue, i.e., $\lambda_{d/2}$, exists only when $d$ is even.
@@
Eigenvalues of $G$ when $d$ is odd. In this case, for $i\in\{1,\ldots,(d-1)/2\}$, the nondegenerate eigenvalues are
@@
Eigenvalues of $G$ when $d$ is even. In this case, the eigenvalues are the same as above, here for $i\in\{1,\ldots,d/2-1\}$, along with the two new eigenvalues $1+(L-1)|N_{d/2,d/2}|$ and $1-|N_{d/2,d/2}|$, which have multiplicities $1$ and L1L-1, respectively.